OpenStudy (anonymous):

x^2 - 7x + 6 = 0 Solve the equation by graphing.

OpenStudy (jannine24):



OpenStudy (anonymous):

This is solving Quadratic Equations by Graphing. We have to show it on a graph. Help please.

OpenStudy (jannine24):

do u need a vertex

OpenStudy (anonymous):

Yes please

OpenStudy (jannine24):

so the formula for vertex is -b/2(a)

OpenStudy (jannine24):

so it would be -(-7))/2(1) right

OpenStudy (anonymous):

I am lost. My teacher wants us to solve the problems with Quadratic Equations by Graphing.

OpenStudy (anonymous):

and we have to show all our work

OpenStudy (jannine24):

r u looking for square root in the quadratic equation

OpenStudy (anonymous):

no

OpenStudy (anonymous):

1. x squared - 7x + 6 = 0

OpenStudy (jannine24):

then what is the equation

OpenStudy (jannine24):

ax2 + bx + c = 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait x2 + bx + c = 0